OpenStudy (anonymous):

Trying to find the circumference of the circle using 22 over seven or 3.14 for pie

OpenStudy (anonymous):

lol pi. do they give you the radius or diameter of this circle?

OpenStudy (anonymous):

yes its 11m

OpenStudy (anonymous):

yes its 11m

OpenStudy (anonymous):

yes its 11m

OpenStudy (anonymous):

yes its 11m

OpenStudy (anonymous):

yes its 11m

OpenStudy (anonymous):

yes its 11m

OpenStudy (anonymous):

what is 11m? the diameter or radius?

OpenStudy (anonymous):

yes its 11m

OpenStudy (anonymous):

yes its 11m

OpenStudy (anonymous):

yes its 11m

OpenStudy (anonymous):

ysorry my computer froz and i pressed it a million times

OpenStudy (anonymous):

haha. is that the diameter or the radius?

OpenStudy (anonymous):

i think its the diameter

OpenStudy (anonymous):

thanks for helping me

OpenStudy (anonymous):

its the radius im sorry

OpenStudy (anonymous):

circumference = 2r*pi = d*pi r=radius and d=diameter